OpenStudy (anonymous):

identify the slope and the y-intercept of the equation. y = negative 1 over 3x + 3 –x + 3y = 9

OpenStudy (anonymous):

\[y = -\frac{1}{3}x + 3\]

OpenStudy (anonymous):

Can you tell me here what is the coefficient of x here??

OpenStudy (anonymous):

is it x=9 @waterineyes

OpenStudy (anonymous):

am so confused

OpenStudy (anonymous):

See if I say: \(y = ax + b\) Then here, \(a\) is the coefficient of x. Similarly what is coefficient of x there?

OpenStudy (anonymous):

am confused

OpenStudy (nubeer):

well which part u r not getting at all ?

OpenStudy (anonymous):

omg i dont even know where to start

OpenStudy (anonymous):

−1 Y= --- X +3 3 is this what your problem looks like?

OpenStudy (anonymous):

Yes @iforgot

OpenStudy (anonymous):

Slope intercept form is this y = mx + b y = x = m= slope b = y intercept

OpenStudy (anonymous):

so what do i have to do?

OpenStudy (anonymous):

well what does you question say i should be asking you :P

OpenStudy (anonymous):

identify the slope and the y-intercept of the equation. @iforgot

OpenStudy (anonymous):

which would be?

OpenStudy (anonymous):

it is already put in slope intercept form you do not have to do anything

OpenStudy (anonymous):